the 1996 Olympic times for Michael Johnson in the 200 m and Donovan Bailey's 100 meter times are shown. Both athletes won their races and won a gold medal. Your task is to explain which athlete had the highest Maximum Instantaneous Velocity during their race. Support for your answer using information in the book, formulas and/or inductive reasoning based on your understanding of these races. You will not be able to calculate the exact maximum velocity, but the material presented in the book and reasoning can give solid support to an answer.

-Donovan Bailey (100m - 9.84s)

average speed 100/9.84 = 10.12 m/s

-Michael Johnson (200m - 19.32s)

average speed 200/19.32 = 10.35 m/s
Answer #1

Maximum instaneous velocity will have maximum avg velocity

Donovan bailey avg velocity 100/9.84 = 10.12 m/s

Michael Johnson avg velocity = 200/19.32 = 10.35 m/s

Avg is (0+ Vmax )/2

So Vmax is higher for michael johnson .

Another reason is continuous acceleration for more time